Tayco Corporation has just paid dividends of $3 per share. The earnings per share for the company was $4. If you believe that the appropriate discount rate is 15% and the long term growth rate in dividends is 6%, and earnings is 6%, the firm’s P/E ratio is

  1. 8.33
  2. 33.33
  3. 44.44
  4. 11.11
  5. None of the above

the answer is A, why?
Answer #1

P/ Ratio = share price/ Earning Per share

The dividend is growing at a rate of 6%

3(1.06)

EPS = 4(1.06)/0.15-1.06

=(3(1.06) 4(1.06}} {0.15 -0.06) = 8.33